OpenStudy (anonymous):

• Provide an example of a trig function and describe how it is transformed from the standard trig function f(x) = sin x, f(x) = cos x, or f(x) = tan x

OpenStudy (welshfella):

an example is the sine and the cosecant

OpenStudy (anonymous):

@welshfella im confused

OpenStudy (anonymous):

Taking the example of Welshfella, the Cosecant function derives from the reciprocal of the Sine function. That is: \[cscx = \frac{ 1 }{ \sin x }\]

OpenStudy (anonymous):

Why is the interval important when calculating the rate of change on a trig function? @Hoslos

OpenStudy (anonymous):

Similarly, the Tangent function derives from the division of sine and cosine: \[tanx = \frac{ sinx }{ \cos x }\]

OpenStudy (anonymous):

That is crucial, as the result of the trig function might give a zero in the denominator, which will make the function undefined. For instance, given the above formula for Tangent, I cannot include the root 90Degrees, because \[\frac{ \sin 90 }{ \cos 90 }=\frac{ 1 }{ 0 }=undefined\] Is that understandable?